OpenStudy (anonymous):

Find an explicit solution of the given initial-value problem. sqrt(1-y^2)dx-sqrt(1-x^2)dy=0 when y(0)=1/2

OpenStudy (anonymous):

i got to the part of y=sin(arcsin(x)+C)

OpenStudy (anonymous):

plugging y0=1/2

OpenStudy (anonymous):

gives 1/2=Sin(C)

OpenStudy (anonymous):

how do i get the constant out by its self?

OpenStudy (anonymous):

is it correct if i jus do arcsin on both sides

OpenStudy (anonymous):

Assuming all your work is right so far, you can isolate C by taking the inverse sine of both sides. So what's \(\arcsin\dfrac{1}{2}\)?

OpenStudy (anonymous):

pi/6

OpenStudy (anonymous):

so can i say my explicit solution is y=sin(arcsin(x)+(pi/6))

OpenStudy (anonymous):

oh nice i totally forgot of wolframalpha today

OpenStudy (anonymous):

Yeah, that's it

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw